OpenStudy (anonymous):

Simplify using positive exponents. m^3n^–6p^0

OpenStudy (anonymous):

\[m^3n ^{-6}p^0 \] is the proper display.

OpenStudy (anonymous):

does it help to know that \(p^0=1\) ?

OpenStudy (astrophysics):

and \[x^{-1} = \frac{ 1 }{ x }\]

OpenStudy (anonymous):

and that \[n^{-6}=\frac{1}{n^6}\]

OpenStudy (anonymous):

so\[\frac{ 2m^3 }{ ?n^6 } \] would be correct to assume?

OpenStudy (astrophysics):

Where'd you get the 2

OpenStudy (anonymous):

ignore the question mark on the bottom

OpenStudy (anonymous):

where did the 2 come from?

OpenStudy (anonymous):

it was a typo it was supposed to be 1

OpenStudy (anonymous):

usually you don't write the 1 either, just go with \[\frac{m^3}{n^6}\]

OpenStudy (anonymous):

So would that be the answer? or is there more to it?